OpenStudy (anonymous):

Evaluate show work 4x+21/y-17 For x = 9 and y = 26. Write your final answer exactly as a fraction in lowest terms

OpenStudy (anonymous):

57/9 divide both by 3 =19/3 so 19/3

OpenStudy (anonymous):

i dont uderstand

OpenStudy (anonymous):

4(9)=36 36+21=57

OpenStudy (anonymous):

thats for x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

26-17=9 for y

OpenStudy (anonymous):

so you have 57 over 9

OpenStudy (anonymous):

divide both of them by 3 making 57 go to 19 and 9divided by 3 is 3 so 19 over 3

OpenStudy (anonymous):

ok so 19/3 do i need to simplify any further

OpenStudy (anonymous):

no because you cant simplify any further

OpenStudy (anonymous):

ok thanx that makes sense i have more problems i will post if u know pls assist...

OpenStudy (anonymous):

ok np!:)

OpenStudy (anonymous):

k